OpenStudy (anonymous):

(x - 8)(x^2 + 7x + 36) = 0 how would i use the quadratic formula to find the complex/imaginary roots?

OpenStudy (anonymous):

On of the roots is x = 8. To find other two, factor x^2 + 7x + 36 and set equal to zero.

OpenStudy (kainui):

You know that (x-8) or (x^2+7x+36) equal 0 since if either one of those terms is equal to zero, then by multiplying the other by 0 will give you 0, makes sense. So x-8=0 easily solves, but the other term uses the quadratic formula, so lets just set that term by itself equal to 0. x^2+7x+36=0 Now you can plug in a, b, and c from that equation into the quadratic formula to find your answers from there.

OpenStudy (anonymous):

x^2+7x+36 = 0 Roots are: (-7 +/- SQRT(49-4*36))/2 => -7 +/- SQRT(-95))/2 => (-7 + i*SQRT(95))/2 OR (-7 - i*SQRT(95))/2

OpenStudy (anonymous):

So wait, The roots are -7 + i*SQRT(95))/2 OR (-7 - i*SQRT(95))/2

OpenStudy (anonymous):

And also 8 as I showed above.

OpenStudy (anonymous):

Wow thank you so much!!!